OpenStudy (anonymous):

Medal 25% of what is 73.5?

OpenStudy (butterflydreamer):

25% of x is 73.5 this means that 0.25x = 73.5 Now we divide both sides by 0.25 to solve for x

OpenStudy (anonymous):

u can also do 73.5 x .25

OpenStudy (anonymous):

18

OpenStudy (anonymous):

yaz

OpenStudy (anonymous):

but that makes no sense

OpenStudy (butterflydreamer):

we're not solving for 25% of 73.5 We want to know what 25% of a NUMBER (x) = 73.5 So this means a QUARTER of a number is 73.5 You need to multiply 73.5 by 4 then to find the original number..

OpenStudy (anonymous):

OH YA

OpenStudy (anonymous):

25% of what is 44.5 = 178 so how can 25% of what is 73.5 be 18

OpenStudy (butterflydreamer):

are you like ignoring what i'm saying lmaonaise.

OpenStudy (anonymous):

ok 294

OpenStudy (anonymous):

73.5 * 4 = 294

OpenStudy (butterflydreamer):

yeah that's the answer lol

OpenStudy (anonymous):

thank you :)

OpenStudy (butterflydreamer):

no problem haha.

OpenStudy (texaschic101):

percent formula : is/of = %/100 25% of what is 73.5 is/of = %/100 73.5/x = 25/100 cross multiply 25x = 73.5(100) 25x = 7350 x = 7350/25 x = 294 this formula works on all percent problems...just make sure you put the numbers in the right place in the formula